VEGAS Team Meeting

Attendees: Amanda, Joe, Marty, Richard, Mark, Randy, Jason, John, Hong

Mute Manager: Mark Whitehead

Agenda:

Whole group:

  1. DONE Status of 1080Mz / 750 MHz builds and MMCM errors - is this really a problem?
    • the problems with the 1000 MHz BOF files
  2. DONE Building new BOFs in Green Bank - status / schedule

GB only

  1. DONE Status of Tsys problem
  2. DONE Ghost frequency issue and plans going forward (software fix or not)
  3. DONE Testing time this week / weekend
  4. DONE Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Meeting notes:

Whole group:

  1. Status of 1080Mz / 750 MHz builds and MMCM errors - is this really a problem?
    • The MMCM failures are an issue - 1080MHz is not the solution
    • the problems with the 1000/1080 MHz BOF files
      • Berkeley is looking at the issue.
      • John proposes for discussion disabling these modes for now, and we should hold up on formal "final" testing until these items are resolved, as we will likely be modifying the designs and/or libraries to fix these problems, and there's no sense in testing things that will have to be retested again later.
    • The 750 BOF still has the ghost frequency
      • Test the 750 MHz version with MMCM - test Tuesday am in GB
  2. Building new BOFs in Green Bank - status / schedule
      • We are doing full regression tests and this is very time consuming. Is there a subset that can be used for a regression/smoke test of a BOF?
        • The MMCM test at a number of rates is a good candidate
        • Resume BOF building to see if it can be sorted, but don't test until the Berkeley-derived changes are identified

GB only

  1. Status of Tsys problem
    • There is a Tsys issue that Anish and possibly Amanda (data in switching mode) have seen in Q-Band and suspected in other bands
      • Does not seem to be cal - addition of frequency switching caused problem
      • What tests should be conducted. Doppler tracking?
      • Additional data analysis by Amanda will begin Wednesday
      • Results to 'gbsapp'
  2. Ghost frequency issue and plans going forward (software fix or not)
    • Simulations continue to test hardware and another data ordering issue may have identified - will cross validate with second investigator
    • Keep up investigations until the BOF issues are resolved and final tests can be run
  3. Use of test time
    • Unscheduled request(s)
      • none
    • Tuesday
      • 08:00 possible carry forward of mixed mode w/new config tool changes
    • Wednesday
      • Available, but Richard will be in CR if there are tests to run.
    • Thursday
      • Placeholder for TSys tests (blanking, doppler, some of Ron's tests) [Kepley]
    • Monday
  4. Status of other “current issues”: https://safe.nrao.edu/wiki/bin/view/CICADA/CICADAGreenBankSpectrometerCurrentIssues

Other

  1. If we see traffic spikes we could plan to meet additionally on Wednesdays
  2. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-07-07
Topic revision: r3 - 2014-07-07, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback